OpenStudy (anonymous):

The average of two students' scores on a math exam is 70. The firs student's score is ten less than twice the score of the second student. What is the score of the first student on the exam? A: 90. Why is this?

OpenStudy (4n1m0s1ty):

Do you know the formula for the average?

OpenStudy (anonymous):

my false logic: 2y-10=70 2y=80 y=40 80-10=70

OpenStudy (anonymous):

the sum of the scores dived by the amount of scores. yes

OpenStudy (anonymous):

divided*

OpenStudy (4n1m0s1ty):

So lets call student 1's score x and student 2's score y. Can you use this to write an algebraic expression?

OpenStudy (anonymous):

thank you waterineyes :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

I have done a little mistake: x = 2y - 10 \[\frac{x + y}{2} = 70 \implies \frac{3y-10}{2} \implies 3y - 10 = 140 \implies 3y = 150\] Now find y from it and then find x..

OpenStudy (anonymous):

oh my... thank you for that. i appreciate that D:

OpenStudy (anonymous):

Welcome dear and sorry...

OpenStudy (anonymous):

oh dont be at all! im just glad im getting help on all these questions D: math is probably my worst subject haha